I'm trying to solve this parabola and I'm kind of stumped
With my equation, which is ... (x+-24)(x+-0) =? and I was told to use the process that we call " foil " but I'm still a little lost.
Thanks :)

Found 2 solutions by richard1234, Theo:
Answer by richard1234(7193)

"Foil" is a mnemonic meaning "first, outer, inner, last." For example, if we have a product in the form (a+b)(c+d), we would multiply the first terms (a, c), followed by the outer terms (a, d), followed by the inner terms (b, c), finally the last terms (b, d). Adding them together, this becomes

ac + ad + bc + bd.

Try that for your problem.

Answer by Theo(13342)

foil is an acronym for first, outer, inner, last.

your factors are (x + 24) * (x + 0)

you have a left factor and a right factor.

x + 24 is the left factor
x + 0 is the right factor

the first terms in each factor are x from the left factor and x from the right factor.

the outer terms in each factor are x from the left factor and 0 from the right factor.

the inner terms in each factor are 24 from the left factor and x from the right factor.

the last terms in each factor are 24 from the left factor and 0 from the right factor.

ussng foil you would

multiply x * x (first terms in each factor)
multiply x * 0 (outer terms in each factor)
multiply 24 * x (nner terms in each factor)
multiply 24 * 0 (last terms in each factor)

you would get x^2 + 0 + 24x + 0 which you would then add together to get:

x^2 + 24x.

that would be your answer.

(x+24) * (x+0) = x^2 + 24x

you can also use the distributive property of arithmetic to get the same result.

(x + 24) * (x + 0) = x * (x + 0) + 24 * (x + 0)

the reslt of this is x^2 + 0 + 24x + 0 which becomes x^2 + 24x

here's a reference on distributive property of multiplication.

http://math.about.com/od/algebra/a/distributive.htm




foil is just another way to accomplish the same thing.